How many moles ofOH- are there in 4 moles ofAl(OH)3
1
Expert's answer
2019-11-15T07:06:51-0500

Solution.

"AL(OH)3 = Al^{3+} + 3OH^{-}"

"n(OH^{-}) = 3 \times n(Al(OH)3)"

"n(OH^-) = 3 \times 4 = 12 \ mol"

Answer:

"n(OH^-) = 3 \times 4 = 12 \ mol"